OpenStudy (anonymous):

Daniel traveled by scooter at an average speed of 28mp/h. Then he traveled by car at an average speed of 44mp/h. In total he traveled 378 miles for 9 hours. How much many minutes did Daniel travel by scooter?

OpenStudy (anonymous):

S = 18/16 = 67.5 min

OpenStudy (anonymous):

could you some me your work?

OpenStudy (anonymous):

Let S is the hours travel by scooter: 28S + 44 ( 9 - S) = 378 miles

OpenStudy (anonymous):

Can you continue from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I thought it was a system of equation problem

OpenStudy (anonymous):

Just multiply it out! It's not as complicated as you thought. Of course it's simple for me to do all the work, but you won't be benefited at all!

OpenStudy (anonymous):

I already knew the answer b4 I posted the question. Could you explain 28S + 44 ( 9 - S) = 378 miles.

OpenStudy (anonymous):

Apply: speed * time = distance

OpenStudy (anonymous):

that explains the 28S but why add the speed of the car?

OpenStudy (anonymous):

Because the distance is total for both scooter and car.

OpenStudy (anonymous):

and the 9-s?